OpenStudy (anonymous):

A construction company orders 2900 new light bulbs for a new office building. If the probability of a light bulb being defective is 1.9%, predict how many of the light bulbs are likely to be defective. Round your answer to the nearest whole number. A. 55 light bulbs B. 69 light bulbs C. 46 light bulbs D. 551 light bulbs

OpenStudy (anonymous):

calculate 1.9% of 2900

OpenStudy (anonymous):

That will be nearly 2.0%

OpenStudy (anonymous):

Am I supposed to divide 1.9% into 2900 to get that? I suck at this ._.

OpenStudy (anonymous):

so, 2900*1.9/100 = 29*1.9 = 55.1 or 55

OpenStudy (anonymous):

so Mam is that okay :)

OpenStudy (anonymous):

Lol I see, thank you [:

OpenStudy (anonymous):

No problem, my pleasure.